# SFO: Piloting VLM Feedback for Offline RL

### v1

@@ 3.1 Sub-Trajectory Filtered Behavior Cloning (SFBC)
A sub-trajectory is retained if $P_{VLM}(\tau_{i})\geq\theta$. To ensure robustness, we employ retrospective filtering, where a sub-trajectory $\tau_{i}$ is only included if the next sub-trajectory $\tau_{i+1}$ also meets the threshold:
$\mathcal{D}_{SFBC}=\{(s_{t},a_{t})\mid\tau_{i}\in\mathcal{D},P_{VLM}(\tau_{i})\geq\theta\text{ and }P_{VLM}(\tau_{i+1})\geq\theta\}.$ (5)
@@
3. Weighted Behavior Cloning: The policy is then trained using the following behavior cloning objective:
$\mathcal{L}_{SFBC}=-\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}_{SFBC}}\left[P_{VLM}(\tau_{t})\log\pi(a_{t}|s_{t})\right].$ (6)
@@ 4.2 Baseline Comparisons
Furthermore, SFBC outperforms VLM+TD3+BC, justifying our approach of treating $P_{VLM}$ as a weight rather than a reward. Conceptually, the weighting scheme can be interpreted as a learned behavioral value function $Q^{\pi_{\text{behavior}}}$, where the weighted loss, $\mathcal{L}_{SFBC}$, implicitly optimizes a policy gradient for the behavior policy that generated the expert sub-trajectories. This provides a principled alternative to direct reward interpretation, avoiding instability in reward propagation.
@@ 4.3 Ablation Studies
  • Markov Prompt Only: Using only the Markovian prompt leads to degraded performance, confirming the importance of leveraging a non-Markovian understanding of trajectory improvement, even in Markovian environments, due to the inability of VLMs to process control data.
@@ A.3 Hyperparameters
We used the default hyperparameters from D3RLpy for all baselines. Additionally, we tested a learning rate of $1e^{-3}$ for AWAC and TD3+BC, as this is the default for BC, but it did not improve results. We found $\theta=0.1$ to be the most effective threshold for filtering.

### v2

@@ 3.1 Sub-Trajectory Filtered Behavior Cloning (SFBC)
A sub-trajectory is retained if $P_{VLM}(\tau_{i})\geq\alpha$. To ensure robustness, we employ retrospective filtering, where a sub-trajectory $\tau_{i}$ is only included if the next sub-trajectory $\tau_{i+1}$ also meets the threshold:
$\mathcal{D}_{SFBC}=\{(s_{t},a_{t})\mid\tau_{i}\in\mathcal{D},P_{VLM}(\tau_{i})\geq\alpha\text{ and }P_{VLM}(\tau_{i+1})\geq\alpha\}.$ (5)
@@
3. Weighted Behavior Cloning: The learned policy, $\pi_{\theta}$, is then trained using the following behavior cloning objective:
$\mathcal{L}_{SFBC}=-\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{D}_{SFBC}}\left[P_{VLM}(\tau_{t})\log\pi_{\theta}(a_{t}|s_{t})\right].$ (6)
@@ 4.2 Baseline Comparisons
Furthermore, SFBC outperforms VLM+TD3+BC, justifying our choice to treat $P_{VLM}$ as a weight rather than a reward. Conceptually, the weighting scheme can be interpreted as a value function $Q^{\pi_{\theta}}$, such that the weighted loss, $\mathcal{L}_{SFBC}$, implicitly optimizes a policy gradient. However, $\mathcal{L}_{SFBC}$ lacks an importance sampling correction, and $P_{VLM}$ is computed from offline data rather than the current policy, making it a better model of $Q^{\pi_{\text{behavior}}}$ (or $Q^{\pi_{\text{expert}}}$ after filtering). SFBC also resembles AWAC (Nair et al.,, 2020), except that the weighting $P_{VLM}$ should likewise be interpreted as approximating $e^{A^{\pi_{\text{expert}}}}$, rather than the advantage of the learned policy $e^{A^{\pi_{\theta}}}$. While SFBC does not fully correspond to a policy gradient, our results show that avoiding a direct reward interpretation—and the instability it introduces in reward propagation—is beneficial in practice.
@@ 4.3 Ablation Studies
  • Markov Prompt Only: Using only the Markov prompt leads to degraded performance, confirming the importance of a non-Markovian understanding of trajectory improvement, even in Markovian environments, due to the inability of VLMs to process control data.
@@ A.3 Hyperparameters
We used the default hyperparameters from D3RLpy for all baselines. Additionally, we tested a learning rate of $1e^{-3}$ for AWAC and TD3+BC, as this is the default for BC, but it did not improve results. We found $\alpha=0.1$ to be the most effective threshold for filtering.